OpenStudy (anonymous):

What is cos(2cos^-1x) written as an algebraic expression in x?

zepdrix (zepdrix):

\(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

\[\cos(2\theta)=2\cos^2(\theta)-1\] should do it

OpenStudy (anonymous):

so \[\cos(2\cos^{-1}(x))=2\cos^2(\cos^{-1}(x))-1\]

OpenStudy (anonymous):

and since \(\cos(\cos^{-1}(x))=x\) you get your algebraic expression